OpenStudy (anonymous):

what is the period of the function? y=2cos pi x

OpenStudy (anonymous):

the period of \(\cos(bx)\) is \(\frac{2\pi}{b}\)

OpenStudy (anonymous):

in your case you have \[\frac{2\pi}{\pi}=2\]